# Identification of cardiovascular diseases through ECG classification using wavelet transformation

### v1

@@ 1 Introduction
Cardiovascular diseases (CVDs) remain the predominant cause of death globally, accounting for an estimated 17.9 million lives each year according to the World Health Organization. The early diagnosis of CVDs is crucial for effective treatment and management. Electrocardiogram (ECG) tests, which record the electrical activity of the heart over a period, play an essential role in the initial screening and diagnosis of heart conditions. Recent advances in methodology and machine learning techniques has prompted health researchers to utilize advance analytics to solve societal and health related issues, providing pathways for experts to tackle issues impacting population. Maleki et al. (2022); Maleki and Khan (2023); Maleki (2024a, b)
However, the manual analysis of ECG data is time-consuming and subject to variability in expert interpretation. As the volume of data generated increases, automated and more accurate methods for ECG analysis become imperative. Machine learning offers promising tools for enhancing the analysis of ECG data, particularly through the application of signal processing techniques.
Among various signal processing methods, wavelet transformation has emerged as a powerful technique for analyzing non-stationary physiological signals. Unlike the Fourier transform, which only provides frequency information, wavelet transforms can capture both time and frequency information, making them particularly suited for the analysis of ECG signals which consist of complex, transient signal components.
@@ 2 Literature Review
The analysis of ECG signals using machine learning techniques has been extensively studied, with various approaches showing promise in enhancing diagnostic accuracy. Early methods predominantly relied on Fourier transforms to analyze the frequency components of ECG signals. However, the Fourier transform’s limitation in capturing time-varying properties of ECG signals often led to inadequate results in real-world scenarios where ECG characteristics fluctuate Smith et al. (2017).
Wavelet transform, introduced as an alternative to Fourier transform, provides a more nuanced analysis by allowing both time and frequency resolution to adapt dynamically Jones et al. (2018). The adaptability of wavelet transforms makes them particularly suited for ECG signals, which contain non-stationary or transient characteristics that are crucial for identifying cardiovascular abnormalities Lee et al. (2019).
@@
Several studies have employed discrete wavelet transforms (DWT) for ECG signal processing. For instance, Addison et al. (2005) demonstrated the potential of DWT in detecting myocardial infarction by extracting detailed features from different ECG signal components Addison (2005). Their work highlighted the ability of wavelet features to improve the sensitivity and specificity of automated diagnostic systems.
Moreover, continuous wavelet transform (CWT) has also been applied to analyze ECG signals for arrhythmia detection. In a comprehensive study by Martins et al. (2018), CWT was used to classify various types of arrhythmias with a high degree of accuracy, using a combination of wavelet-based features and machine learning classifiers Martins et al. (2018).
@@
Recent advances have integrated wavelet transformation with deep learning techniques to further enhance the classification accuracy. Zhang et al. (2020) developed a convolutional neural network model that uses wavelet-transformed ECG signals to automatically detect arrhythmias, achieving state-of-the-art performance Zhang et al. (2020). This integration of wavelet analysis and deep learning underscores the potential for significant improvements in automated ECG analysis systems.
However, despite these advancements, the application of wavelet transforms in ECG classification still presents challenges, particularly in handling large datasets and optimizing feature selection for improved classifier performance. This study aims to address these challenges by exploring both the discrete and continuous wavelet transformations to evaluate their efficacy in feature extraction and subsequent classification accuracy in a large-scale ECG dataset.
3 Materials and Methods
3.1 Data Source
@@
The study utilizes the MIT-BIH Arrhythmia Database, which comprises 48 half-hour excerpts of two-channel ambulatory ECG recordings from 47 subjects studied at the Beth Israel Hospital between 1975 and 1979 Goldberger et al. (2000). The dataset is a well-established resource in the field of cardiovascular disease research and provides a diverse range of ECG signal patterns for analysis. A processed and segmented version of this dataset, prepared for a Kaggle challenge by Kachuee et al. (2018), is used, which includes 109,446 samples categorized into five classes Kachuee et al. (2018). The samples are pre-divided into 87,554 training and 21,892 testing samples, with this study using a 70/30 split of the training dataset for model training and validation.
@@ 3.2 Wavelet Transformation
A typical ECG signal consists of a number of underlying oscillatory frequency components superimposed by a noise component. A general signal is a superposition of special scales at different locations in time. Unlike Fourier transform that decomposes a signal into a number of scales using either time or frequency, wavelet transform decomposes a signal into a set of basic functions called wavelets. While Fourier transform uses a linear combination of infinitely-long sine-waves to represent a signal, wavelet transform uses a series of localized wavelets each with different scale, which allows for obtaining time information as well as the frequency information (Figure 1).
@@
The continuous wavelet transform (CWT) is described as follow Valens (1999):
@@
which shows how a signal function $f(t)$ can be decomposed into a set of wavelets $\Psi_{s,\tau}(t)$ based on scale and position parameters, a and b respectively. The wavelet family members for a signal are generated from a single basic wavelet (mother wavelet) by changing the scale (compressing or stretching) and shifting the position as formulated below:
@@
To focus on a more manageable set of wavelets produced from an infinite number of wavelets in the continuous wavelet transformation and increase the efficiency, discrete wavelet transformation (DWT) is introduced. Unlike continuous, these wavelets can only be scaled and shifted in discrete steps.
@@
Figure 4: Schematic of the classification approach, in which the outputs of DWT function are used to extract statistical features and incorporate them as inputs into the classifier
The characteristic behavior of a signal is mirrored in the frequency sub-bands. By collecting statistical features from each of the frequency sub-bands and using them as inputs to train a classifier, we can characterize a signal and predict the class or disease it is representing. Selection of statistical features is highly dependent on the type of signals. In this project, the following 8 features will be investigated: mean, median, standard deviation, variance, root mean square value, zero crossings rate, mean crossing rate, and entropy inspired by Taspinar (2018).
After extracting several statistical features for each sub-band of each signal, they will all be combined into a single dataset and used as inputs for a classifier. Figure 4 describes the approach to convert an ECG signal into features and used them as inputs for classification.
Both Continuous Wavelet Transform (CWT) and Discrete Wavelet Transform (DWT) are applied. The CWT offers a continuous view of the signal across all scales and translations, which is beneficial for detailed analysis. However, to focus on a more manageable set of wavelets and increase computational efficiency, the DWT is predominantly used. The DWT splits the signal into a series of frequency sub-bands using high-pass and low-pass filters, which return detail and approximation coefficients, respectively Mallat (1989). This process is repeated on the approximation coefficients from the previous level of DWT, effectively decomposing the signal into multiple levels, each corresponding to different frequency bands.
@@ 3.3 Feature Extraction
From each level of DWT decomposition, several statistical features are extracted to represent the characteristics of the signal in each frequency band. These features include the mean, median, standard deviation, variance, root mean square value, zero-crossing rate, mean crossing rate, and entropy Taspinar (2018). The extraction of these features is inspired by the ability of such metrics to capture essential characteristics of the signal relevant to differentiating between normal and abnormal ECG patterns. All extracted features from the sub-bands are then combined into a comprehensive feature vector for each sample.
@@ 3.4 Classification Models
A range of machine learning classifiers are evaluated for their ability to classify ECG signals based on the extracted features. These classifiers include K-Nearest Neighbors, Support Vector Machines with linear and radial basis function kernels, Decision Trees, Random Forests, Multi-layer Perceptrons, AdaBoost, Gaussian Naive Bayes, and Gradient Boosting. Each model’s performance is assessed based on accuracy, precision, recall, and F1-score, with a focus on optimizing hyperparameters to improve prediction accuracy and model robustness Pedregosa et al. (2011).
@@ 3.5 Implementation Details
The implementation of the transformation, feature extraction, and classification processes is carried out using Python, leveraging libraries such as PyWavelets for wavelet transformations and scikit-learn for machine learning modeling. The computational work is managed within a Jupyter Notebook environment to ensure reproducibility and facilitate documentation of the code and results.
@@ 4 Results
The classification of ECG signals involved processing signals that represent both normal and various abnormal heart conditions, including different arrhythmias and myocardial infarction. These signals were preprocessed and segmented, with each segment representing a single heartbeat, as depicted in Figure 5.
Figure 5: Example of five types of heartbeats signals corresponding to five classes of arrhythmias and myocardial infarction
@@ 4.1 Wavelet Choice and Decomposition
For signal decomposition, the "sym5" wavelet was chosen due to its nearly symmetric properties and effectiveness in decomposing non-stationary signals like ECG. This choice is crucial as it impacts the quality of the feature extraction process. The sym5 wavelet, a member of the Daubechies family, is known for its balance between time and frequency localization. Figure 6 illustrates the typical shape of the sym5 wavelet used in this study.
Figure 6: Typical shape of sym5 wavelet used to decompose the ECG signals
@@ 4.2 Feature Extraction and Dataset Preparation
Using the discrete wavelet transform (DWT), each ECG signal was decomposed into five levels. From each level of decomposition, eight statistical features were extracted, including mean, median, standard deviation, variance, root mean square value, zero-crossing rate, mean crossing rate, and entropy. This resulted in a comprehensive dataset with 40 features per sample, derived from different frequency sub-bands of each decomposed signal. The dataset was then split into a training set (70%) and a testing set (30%).
@@ 4.3 Classifier Performance and Evaluation
The performance of various classifiers was evaluated, as summarized in the table below, which lists the classifiers along with their hyperparameters, and their respective training and testing accuracies:
@@
RBF SVC gamma=2, C=1 0.99 0.83
Decision Tree max_depth=20 0.99 0.93
@@
Multi-Layer Perceptron hidden_layer_sizes = (50,100), alpha = 0.01, max_iter=1000, activation = ‘tanh’, solver=’adam’ 0.97 0.95
AdaBoost N/A 0.83 0.82
GaussianNB N/A 0.67 0.66

### v2

@@ 1 Introduction
Cardiovascular diseases (CVDs) remain the predominant cause of death globally, accounting for an estimated 17.9 million lives each year according to the World Health Organization World Health Organization (2021). The early diagnosis of CVDs is crucial for effective treatment and management. Electrocardiogram (ECG) tests, which record the electrical activity of the heart over a period, play an essential role in the initial screening and diagnosis of heart conditions.
However, the manual analysis of ECG data is time-consuming and subject to variability in expert interpretation. As the volume of data generated increases, automated and more accurate methods for ECG analysis become imperative. Recent advances in methodology and machine learning techniques has prompted health researchers to utilize advanced analytics to solve societal and health-related issues, providing pathways for experts to tackle issues impacting population. Machine learning offers promising tools for enhancing the analysis of ECG data, particularly through the application of signal processing techniques.
Among various signal processing methods, wavelet transformation has emerged as a powerful technique for analyzing non-stationary physiological signals Addison (2005). Unlike the Fourier transform, which only provides frequency information, wavelet transforms can capture both time and frequency information, making them particularly suited for the analysis of ECG signals which consist of complex, transient signal components.
@@ 1.1 Literature Review
The analysis of ECG signals using machine learning techniques has been extensively studied, with various approaches showing promise in enhancing diagnostic accuracy. Early methods predominantly relied on Fourier transforms to analyze the frequency components of ECG signals Addison (2005); Martins and Jones (2018). However, the Fourier transform’s limitation in capturing time-varying properties of ECG signals often led to inadequate results in real-world scenarios where ECG characteristics fluctuate He and Wu (2019). Wavelet transform, introduced as an alternative to Fourier transform, provides a more nuanced analysis by allowing both time and frequency resolution to adapt dynamically. The adaptability of wavelet transforms makes them particularly suited for ECG signals, which contain non-stationary or transient characteristics that are crucial for identifying cardiovascular abnormalities.
@@
Several studies have employed discrete wavelet transforms (DWT) for ECG signal processing. For instance, Addison et al. (2005) demonstrated the potential of DWT in detecting myocardial infarction by extracting detailed features from different ECG signal components Addison (2005). Their work highlighted the ability of wavelet features to improve the sensitivity and specificity of automated diagnostic systems. Moreover, continuous wavelet transform (CWT) has also been applied to analyze ECG signals for arrhythmia detection. In a comprehensive study by Martins et al. (2018), CWT was used to classify various types of arrhythmias with a high degree of accuracy, using a combination of wavelet-based features and machine learning classifiers Martins et al. (2018).
@@
Recent advances have integrated wavelet transformation with deep learning techniques to further enhance the classification accuracy. Zhang et al. (2020) developed a convolutional neural network model that uses wavelet-transformed ECG signals to automatically detect arrhythmias, achieving state-of-the-art performance Zhang et al. (2020). This integration of wavelet analysis and deep learning underscores the potential for significant improvements in automated ECG analysis systems. However, despite these advancements, the application of wavelet transforms in ECG classification still presents challenges, particularly in handling large datasets and optimizing feature selection for improved classifier performance. This study aims to address these challenges by exploring both the discrete and continuous wavelet transformations to evaluate their efficacy in feature extraction and subsequent classification accuracy in a large-scale ECG dataset.
2 Material and Methods
2.1 Data Sources
@@
The study utilizes the MIT-BIH Arrhythmia Database PhysioNet (2024), which comprises 48 half-hour excerpts of two-channel ambulatory ECG recordings from 47 subjects studied at the Beth Israel Hospital between 1975 and 1979. The dataset is a well-established resource in the field of cardiovascular disease research and provides a diverse range of ECG signal patterns for analysis. A processed and segmented version of this dataset, prepared for a Kaggle challenge by Kachuee et al., is used, which includes 109,446 samples categorized into five classes. The samples are pre-divided into 87,554 training and 21,892 testing samples, with this study using a 70/30 split of the training dataset for model training and validation Kachuee et al. (2018).
@@ 2.2 Wavelet Transformation
A typical ECG signal consists of a number of underlying oscillatory frequency components superimposed by a noise component. A general signal is a superposition of special scales at different locations in time. Unlike Fourier transform that decomposes a signal into a number of scales using either time or frequency, wavelet transform decomposes a signal into a set of basic functions called wavelets (Figure 1). While Fourier transform uses a linear combination of infinitely-long sine-waves to represent a signal, wavelet transform uses a series of localized wavelets each with different scale, which allows for obtaining time information as well as the frequency information.
@@
The CWT is described as follows:
@@
which shows how a signal function $f(t)$ can be decomposed into a set of wavelets $\Psi_{a,b}(t)$ based on scale and position parameters, $a$ and $b$ respectively. The wavelet family members for a signal are generated from a single basic wavelet (mother wavelet) by changing the scale (compressing or stretching) and shifting the position as formulated in Eq. 2:
@@
Figure 4: Schematic of the classification approach, in which the outputs of DWT function are used to extract statistical features and incorporate them as inputs into the classifier. The original schema is reported in Taspinar (2018) article, and utilized here with some modifications.
To focus on a more manageable set of wavelets produced from an infinite number of wavelets in the continuous wavelet transformation and increase the efficiency, DWT is introduced. Unlike continuous, these wavelets can only be scaled and shifted in discrete steps.
@@ 2.3 Feature Extraction
From each level of DWT decomposition, several statistical features are extracted to represent the characteristics of the signal in each frequency band. These features include the mean, median, standard deviation, variance, root mean square value, zero-crossing rate, mean crossing rate, and entropy. The extraction of these features is inspired by the ability of such metrics to capture essential characteristics of the signal relevant to differentiating between normal and abnormal ECG patterns. All extracted features from the sub-bands are then combined into a comprehensive feature vector for each sample.
@@ 2.4 Classification Models
A range of machine learning classifiers are evaluated for their ability to classify ECG signals based on the extracted features. These classifiers include K-Nearest Neighbors, Support Vector Machines (SVC) with linear and Radial Basis Function (RBF) kernels, Decision Trees, Random Forests, Multi-layer Perceptrons, AdaBoost, Gaussian Naive Bayes (NB), and Gradient Boosting. Each model’s performance is assessed based on accuracy, precision, recall, and F1-score, with a focus on optimizing hyperparameters to improve prediction accuracy and model robustness.
@@ 2.5 Implementation Details
The implementation of the transformation, feature extraction, and classification processes is carried out using Python, leveraging libraries such as PyWavelets for wavelet transformations and scikit-learn for machine learning modeling. The computational work is managed within a Jupyter Notebook environment to ensure reproducibility and facilitate documentation of the code and results. The analysis data and code are provided in the following GitHub repository: GitHub - Wavelet Signaling
@@ 3 Results
The classification of ECG signals involved processing signals that represent both normal and various abnormal heart conditions, including different types of arrhythmias and myocardial infarction. These signals were preprocessed and segmented, with each segment representing a single heartbeat, as depicted in Figure LABEL:figure5.
@@ 3.1 Wavelet Choice and Decomposition
For signal decomposition, the "sym5" wavelet was chosen due to its nearly symmetric properties and effectiveness in decomposing non-stationary signals like ECG. This choice is crucial as it impacts the quality of the feature extraction process. The sym5 wavelet, a member of the Daubechies family, is known for its balance between time and frequency localization Mallat (1989). Figure 5 illustrates the typical shape of the sym5 wavelet used in this study.
@@ 3.2 Feature Extraction and Dataset Preparation
Using the DWT, each ECG signal was decomposed into five levels. From each level of decomposition, eight statistical features were extracted, including mean, median, standard deviation, variance, root mean square value, zero-crossing rate, mean crossing rate, and entropy. This resulted in a comprehensive dataset with 40 features per sample, derived from different frequency sub-bands of each decomposed signal. The dataset was then split into a training set (70%) and a testing set (30%).
@@ 3.3 Classifier Performance and Evaluation
The performance of various classifiers was evaluated, as summarized in the Table 1 and Figure 6, which lists the classifiers along with their hyperparameters, and their respective training and testing accuracies:
@@
RBF SVC gamma=2, C=1 1.00 0.83
Decision Tree max_depth=20 0.99 0.94
@@
Multi-Layer Perceptron hidden_layer_sizes = (50,100), alpha = 0.01, max_iter=1000, activation = ‘tanh’, solver=’adam’ 0.96 0.95
AdaBoost N/A 0.85 0.85
GaussianNB N/A 0.66 0.66
@@
Figure 5: Typical shape of sym5 wavelet used to decompose the ECG signals.